OpenStudy (anonymous):

One number is 2 times a first number. A third number is 100 more than the first number. If the sum of the three numbers is 288, find the numbers. The three numbers are?

hartnn (hartnn):

let the 1st number be 'x' "One number is 2 times a first number" so what about that number ? 2x or x+2 ?

OpenStudy (anonymous):

let x,y,z be the 3 numbers y = 2x z = 100 + x x + y + z = 288

OpenStudy (anonymous):

move all the variables to one side, then solve using gaussian algorithm

OpenStudy (anonymous):

ill help you start -2x +y + 0z + 0= 0 x +0y -z + 100 = 0 x + y + z +0 = 0

OpenStudy (anonymous):

I think you guys are making this more complicated than it has to be. I am really confused. "gaussian algorithm"??

hartnn (hartnn):

let the 1st number be 'x' "One number is 2 times a first number" so what about that number ? 2x or x+2 ?

OpenStudy (anonymous):

It's x+2

OpenStudy (anonymous):

you have a system of 3 equations and 3 variables, substitute and subtract until you solve each 3 it's all tedious work from there all the heavy lifting is done

OpenStudy (anonymous):

y = 2x z = 100 + x x + y + z = 288 first one is already done for you, sub 2x for every y you see. substitute 100 + x for every z the last one becomes x + 2x + x + 100 = 288. You can solve that right?

hartnn (hartnn):

2 times implies multiplication by 2 so it must be 2x ok?

OpenStudy (anonymous):

47, 94, 147 i did it my head already

OpenStudy (anonymous):

for the solution just reread each response

OpenStudy (anonymous):

I don't want the answer. I want to know how to do it. Okay, so what do I do after I get 2x hartnn?

hartnn (hartnn):

ok so the other number is 2x now "A third number is 100 more than the first number." so, what can you say about the 3rd number ? x+100, x-100, 100x ...or anything else ?

OpenStudy (anonymous):

x+100

hartnn (hartnn):

correct!

OpenStudy (anonymous):

So, "2x + x + 100" is what we got so far

hartnn (hartnn):

the sum of these 3 will be x+2x+100+x right ? yes! and thats given as 288 so x+2x+100+x=288 can u find x from here ?

hartnn (hartnn):

don't forget the 1st number 'x' :P

OpenStudy (anonymous):

Okay, so 4x+100=288 so this equals= 47

OpenStudy (anonymous):

That's our first number. 47.

hartnn (hartnn):

nice work! correct :)

hartnn (hartnn):

what about the 2nd ?

OpenStudy (anonymous):

hmm so "2 twice the first number"= 94

hartnn (hartnn):

correct again 2nd number = 94 what about 3rd ?

OpenStudy (anonymous):

ok, use the answer to check your work after. The trick in these kinds of questions is to assign each number a variable and for each sentence given, write out a statement in terms of the variables For example: second number is 2 times the first becomes y = 2x and so on

OpenStudy (anonymous):

the once you have your equations you use your substitution and addition/subtraction to solve for each variable

OpenStudy (anonymous):

"a third number is 100 more than the first number"= 100+47= 147 so, 47, 94, 147.

hartnn (hartnn):

you got it :)

OpenStudy (anonymous):

Finally! lol thanks so much guys! Special thanks to you @hartnn Thanks so much! :))

hartnn (hartnn):

welcome so much ^_^